OpenStudy (anonymous):

SOMEONE PLEASE HELP WITH FRACTIONS! 1 5/6 x 3 3/5

OpenStudy (anonymous):

@johnweldon1993 Please show work

OpenStudy (juicstice):

so multiply the whole number with the denominator and then add that to the numerator and then just multiply the top with the top and the bottom with the bottom

OpenStudy (johnweldon1993):

Right...you have 2 mixed numbers...you will want to change those to improper fractions first... \[\large 1\frac{5}{6} \times 3\frac{3}{5}\] To change them to improper fraction...take the whole number....multiply it by the denominator (bottom of the fraction) and then add the numerator (top part of the fraction) to the result...then place that number over the original denominator (bottom) Ill do the first one.... 1 and 5/6 multiply 1 by 6.....= 6 and add the numerator ...6 + 5 = 11 Now place it over the original denominator...makes it 11/6 make sense? now you do the second one :)